A 90 kg skier at the top of a 400 m icy hill slides down, through a 20 m patch of snow (uk= 0.25) and hits a 20 m long jump rb) (5 points) How high will he get of the ground (max height (m)) ?
Answer #1

toon GAL 20m, 115 h = 20 sinas [h= 14:14 m] from every conservation! – (400 - h) mg + Of = 1 hva wo =- Mmg x2 (400 – 14.14)xb) hex Vy 2 = 29 (61.09) 2 c 2x9.8 (hexa hrax = 190.4m Answey CS Scanned with cam scanner